# Text2Traffic: A Text-to-Image Generation and Editing Method for Traffic Scenes

### v1

@@ 1 INTRODUCTION
In recent years, Diffusion Models have driven the rapid advancement of Text-to-Image Generation and Image Editing technologies, thanks to their exceptional generative capabilities and ability to model complex data distributions. [1] These technologies are not only widely applied in fields such as artistic creation, virtual reality, and content generation but are also beginning to demonstrate potential in specialized vertical domains—particularly in Intelligent Transportation Systems. By generating or editing traffic scene images based on natural language instructions, these technologies can provide rich, controllable, and diverse visual data for tasks such as traffic monitoring, autonomous driving simulation, and road safety analysis. This, in turn, helps alleviate bottlenecks associated with the high cost of real data collection, difficulties in annotation, and limited scene coverage.
@@ 2.1 Diffusion Models
In the field of text-to-image generation and image editing, diffusion models have become the mainstream paradigm. Stable Diffusion(SD) [2] effectively alleviates the computational bottleneck encountered in training within high-dimensional pixel spaces by transferring the diffusion process to a latent space and introducing text-conditioned control, while achieving high-quality text-to-image generation. The subsequent SDXL [3] adopts a ”Base + Refiner” two-stage architecture, where the Base model generates a low-resolution image, and the Refiner model enhances details in the high-resolution latent space, significantly improving the fidelity and semantic consistency of the generated images. To further enhance spatial control over the generation process, ControlNet [4] was proposed. By introducing pluggable auxiliary networks, it supports precise guidance from various spatial conditions, such as edge maps, depth maps, and pose maps without compromising the original pre-trained model, greatly expanding the application boundaries of image editing and controlled generation.
Recently, Diffusion Transformer(DiT) [5] has integrated the transformer architecture into diffusion models, replacing the convolutional structures in traditional U-Nets with self-attention mechanisms to better model long-range dependencies and global semantic relationships. DiT has great scalability and generation quality, making it the basis for the next generation of text-to-image models. Building on this foundation, models such as FLUX [6], SD3 [7], Lumina Image 2.0 [8], HunyuanImage 3.0 [9], SeedDream 4.0 [10] and Qwen-Image [11] have expanded based on the DiT architecture, achieving significant advancements in generation quality, multilingual understanding, and multimodal alignment, progressively moving towards a unified multimodal generation representation. However, their effectiveness in text to image and image editing tasks in traffic scenes is not commendable.
Concurrently, multimodal large language models (MLLMs) like Qwen2.5-VL [12] and InternVL3 [13] are widely used for the automatic generation of high-quality image descriptions. These models can generate semantically rich and fine-grained textual annotations, providing more precise supervision signals for text-to-image models, further enhancing the coherence and diversity between generated images and text prompts.
@@ 2.2 Image Generation and Editing in Traffic Scenes
In recent years, text-to-image generation and image editing for traffic scenes have gained increasing attention. Text2Street [14], for instance, utilizes a lane-aware road topology generator to control road layouts in a text-to-map manner, enabling the creation of street scene images from textual descriptions. BEVGen [15] combines an innovative cross-view transformation with spatial attention design, learning the relationship between camera and map views to synthesize realistic and spatially consistent surrounding images that align with the bird’s-eye view (BEV) layout of traffic scenes. DriveDreamer [16] focuses on generating high-quality, controllable driving videos that reflect real-world traffic structures. In the realm of image editing, LightDiff [17] uses a multi-condition controlled diffusion model, incorporating depth maps, RGB images, and text captions to effectively illuminate dark scenes, aiming to enhance low-light image quality for autonomous driving applications. In this paper, we integrate image generation and editing from both vehicle-side and roadside camera perspectives in traffic scenes. This method supports the creation of a wider range of traffic elements and scenes, enhancing their diversity and fidelity. As a result, it provides more valuable data for applications such as autonomous driving, V2X, and beyond.
@@ 3.1 A Multi-Level Captioning Framework
To address these challenges and construct a well-aligned text-to-image traffic dataset, we developed a refined data preparation pipeline shown in Fig. 3. This process aims to generate high-quality, multi-modal annotations necessary for robust model training in traffic domains. Specifically, the pipeline begins with the acquisition of multi-resolution and multi-view traffic images, captured from both roadside and vehicle-side perspectives to ensure holistic geometric consistency and semantic completeness of the scene. A fundamental and distinctive step in our methodology involves processing this domain-specific data with an open-set detection model, such as GroundingDINO [18]. Following the initial detection, we apply Non-Maximum Suppression (NMS) to eliminate redundant and overlapping bounding boxes, ensuring that each detected object is uniquely and accurately localized. Furthermore, to enhance the quality and reliability of the detected objects, we implement a size-based filtering strategy that removes exceptionally small objects lacking distinctive features. This filtering process is essential for maintaining object integrity and preventing the inclusion of ambiguous or noisy visual elements that could hinder model performance.
@@ 3.2 Restoration-based Supervised Fine-Tuning
Fig. 4 illustrates the architecture of our proposed R-SFT framework, a novel training paradigm designed to unify and enhance image generation and editing capabilities within a single model. The framework is structured into three core, interconnected components: (a) Random Mask, (b) Fine-Tuning, and (c) Prompt Alignment, which collectively guide the model from data preparation to high-fidelity output generation. The process initiates with the random mask module (a), which preprocesses the original training images by applying variable-shaped and sized occlusions. This deliberate corruption transforms the task into the restoration problem, forcing the model to learn robust contextual reasoning and inpainting skills essential for both conditional generation and editing. The fine-tuning component (b) forms the core of the training pipeline, implemented as a two-stage progressive learning strategy. In Stage 1, the model is trained to reconstruct images from large-area masks, focusing on recovering broad scene context and global consistency. Stage 2 refines these capabilities using smaller, object-level masks, honing the model’s ability to perform detailed, precise restoration. Both stages leverage a single-stream diffusion transformer  [8] as a backbone, which is efficiently optimized using a combination of LoRA [19] and FFN fine-tuning. The details of the DiT architecture are shown in Fig. 5. In the context of diffusion transformer models, a notable advantage of the single-stream architecture lies in its capacity to seamlessly incorporate supplementary conditional inputs. Unlike alternative approaches that necessitate the introduction of additional module branches, such as adapters or ControlNet [4], the single-stream DiT can directly integrate conditional information without the need for such structural modifications. To leverage this characteristic for the inpainting task, we introduce a dedicated conditional embedder. This embedder is specifically designed to process the inpainting image, denoted as $I_{m}$, obtained by element-wise multiplication of the original image $I$ and the mask $m$. The output of the conditional embedder is then concatenated with noisy latent representations and text tokens. Subsequently, a self-attention mechanism is employed to facilitate multi-modal interaction among these concatenated features. This approach enables the model to effectively combine visual information from the inpainting region, latent representations of the corrupted image, and semantic information from the text prompt, thereby enhancing the overall inpainting performance.
@@
Crucially, the Prompt Alignment module (c) operates during inference to ensure semantic consistency. It utilizes a powerful pretrained Large-Language Model (Qwen-3B [20]) to interpret the user’s textual prompt and align it to the caption distribution. This mechanism provides explicit semantic guidance, enabling the model to perform prompt-aware restoration and object edits. By integrating a restoration-based objective with a structured fine-tuning schedule and a dedicated prompt-alignment mechanism, the R-SFT framework effectively equips a unified model with powerful and versatile image manipulation capabilities.
@@ 3.3 Training objective
We employ flow matching to optimize the parameters of our model. In contrast to Denoising Diffusion Probabilistic Models (DDPM) [1], flow matching performs the forward process by linearly interpolating between noise and data along a straight trajectory. At any given timestep $t$, the noisy sample $x_{t}$ is defined as: $x_{t}=t\cdot x+(1-t)\cdot\epsilon$, where x represents the original data and $\epsilon\sim\mathcal{N}(0,I)$ denotes Gaussian noise. The model is trained to directly regress the target velocity vector $v$, conditioned on the noisy input $x_{t}$, the timestep $t$, and any auxiliary conditioning information $c$. The flow matching loss:
@@
Furthermore, to enhance the model’s focus on the masked regions that require inpainting or editing and to mitigate the common issue of object-level generation, we introduced a weighted loss:
$w_{i,j}=\left\{\begin{array}[]{ll}1&\text{if }\mathbf{I}_{i,j}=\mathbf{I}_{i,j}^{\prime}\\ \frac{1}{\|\mathbf{I}-\mathbf{I}^{\prime}\|^{2}}&\text{if }\mathbf{I}_{i,j}\neq\mathbf{I}_{i,j}^{\prime}\end{array}\right$.
@@ 4.1 Experiment Details
Our dataset consists of three components: open-source datasets (BDD100K [21], TT100K [22], CUHK-SYSU [23]), an internal proprietary dataset, and an internet dataset. The training set comprises approximately 2.4M text-image pairs, including 2M pairs with detailed global image descriptions and 0.4M pairs with local object-level descriptions. The validation set is carefully curated, consisting of 100 samples. This data encompasses a wide variety of scenes in the traffic domain, including different perspectives, elements, weather conditions, and lighting variations. Our model is trained on 8xA800 GPUs using the AdamW [24] optimizer, with an initial learning rate of 1e-4 and a global batch size of 32. The training supports a dynamic resolution strategy and is conducted in two stages. In the first stage, we train for 16 epochs using global image descriptions to capture the high-level semantic structure of traffic scenes. In the second stage, we fine-tune for 5 epochs, using local object descriptions to refine details and enhance local realism. Evaluation metrics such as FID [25], SSIM, and PSNR are used to assess the model’s convergence and generation quality.
@@ 4.3 Quantative Analysis
We evaluated our two-stage model on the validation set, as shown in Tab. 1. In Stage 1, we pre-trained the model using large-scale image-text pairs, achieving FID = 9.8, SSIM = 0.77, and PSNR = 21.06. In Stage 2, we fine-tuned the model with fine-grained object descriptions, resulting in improved performance: FID decreased to 9.6, while SSIM and PSNR increased to 0.81 and 24.52, respectively. An ablation study confirmed that our masked region-weighted loss function further enhanced PSNR from 23.6 to 24.52, demonstrating that focusing loss weighting on object areas can improve detail fidelity and semantic accuracy.

### v3

@@ 1 INTRODUCTION
In recent years, Diffusion Models have driven the rapid advancement of Text-to-Image Generation and Image Editing technologies, thanks to their exceptional generative capabilities and ability to model complex data distributions. [ho2020denoising] These technologies are not only widely applied in fields such as artistic creation, virtual reality, and content generation but are also beginning to demonstrate potential in specialized vertical domains—particularly in Intelligent Transportation Systems. By generating or editing traffic scene images based on natural language instructions, these technologies can provide rich, controllable, and diverse visual data for tasks such as traffic monitoring, autonomous driving simulation, and road safety analysis. This, in turn, helps alleviate bottlenecks associated with the high cost of real data collection, difficulties in annotation, and limited scene coverage.
@@ 2.1 Diffusion Models
In the field of text-to-image generation and image editing, diffusion models have become the mainstream paradigm. Stable Diffusion(SD) [rombach2022high] effectively alleviates the computational bottleneck encountered in training within high-dimensional pixel spaces by transferring the diffusion process to a latent space and introducing text-conditioned control, while achieving high-quality text-to-image generation. The subsequent SDXL [podell2023sdxl] adopts a ”Base + Refiner” two-stage architecture, where the Base model generates a low-resolution image, and the Refiner model enhances details in the high-resolution latent space, significantly improving the fidelity and semantic consistency of the generated images. To further enhance spatial control over the generation process, ControlNet [zhang2023adding] was proposed. By introducing pluggable auxiliary networks, it supports precise guidance from various spatial conditions, such as edge maps, depth maps, and pose maps without compromising the original pre-trained model, greatly expanding the application boundaries of image editing and controlled generation.
Recently, Diffusion Transformer(DiT) [peebles2023scalable] has integrated the transformer architecture into diffusion models, replacing the convolutional structures in traditional U-Nets with self-attention mechanisms to better model long-range dependencies and global semantic relationships. DiT has great scalability and generation quality, making it the basis for the next generation of text-to-image models. Building on this foundation, models such as FLUX [flux], SD3 [esser2024scaling], Lumina Image 2.0 [qin2025lumina], HunyuanImage 3.0 [cao2025hunyuanimage], SeedDream 4.0 [seedream2025seedream] and Qwen-Image [wu2025qwen] have expanded based on the DiT architecture, achieving significant advancements in generation quality, multilingual understanding, and multimodal alignment, progressively moving towards a unified multimodal generation representation. However, their effectiveness in text to image and image editing tasks in traffic scenes is not commendable.
Concurrently, multimodal large language models (MLLMs) like Qwen2.5-VL [bai2025qwen2] and InternVL3 [zhu2025internvl3] are widely used for the automatic generation of high-quality image descriptions. These models can generate semantically rich and fine-grained textual annotations, providing more precise supervision signals for text-to-image models, further enhancing the coherence and diversity between generated images and text prompts.
@@ 2.2 Image Generation and Editing in Traffic Scenes
In recent years, text-to-image generation and image editing for traffic scenes have gained increasing attention. Text2Street [gu2025text2street], for instance, utilizes a lane-aware road topology generator to control road layouts in a text-to-map manner, enabling the creation of street scene images from textual descriptions. BEVGen [swerdlow2024street] combines an innovative cross-view transformation with spatial attention design, learning the relationship between camera and map views to synthesize realistic and spatially consistent surrounding images that align with the bird’s-eye view (BEV) layout of traffic scenes. DriveDreamer [wang2024drivedreamer] focuses on generating high-quality, controllable driving videos that reflect real-world traffic structures. In the realm of image editing, LightDiff [chen2024lightdiff] uses a multi-condition controlled diffusion model, incorporating depth maps, RGB images, and text captions to effectively illuminate dark scenes, aiming to enhance low-light image quality for autonomous driving applications. However, despite these approaches introducing various generation conditions, they are unable to support text-based multi-view traffic scene image generation and editing, thereby limiting the generation of specific scene data. In this paper, we integrate image generation and editing from both vehicle-side and roadside camera perspectives in traffic scenes. This method supports the creation of a wider range of traffic elements and scenes, enhancing their diversity and fidelity. As a result, it provides more valuable data for applications such as autonomous driving, V2X, and beyond.
@@ 3.1 A Multi-Level Captioning Framework
To address these challenges and construct a well-aligned text-to-image traffic dataset, we developed a refined data preparation pipeline shown in Fig. 3. This process aims to generate high-quality, multi-modal annotations necessary for robust model training in traffic domains. Specifically, the pipeline begins with the acquisition of multi-resolution and multi-view traffic images, captured from both roadside and vehicle-side perspectives to ensure holistic geometric consistency and semantic completeness of the scene. A fundamental and distinctive step in our methodology involves processing this domain-specific data with an open-set detection model, such as GroundingDINO [liu2024grounding]. Following the initial detection, we apply Non-Maximum Suppression (NMS) to eliminate redundant and overlapping bounding boxes, ensuring that each detected object is uniquely and accurately localized. Furthermore, to enhance the quality and reliability of the detected objects, we implement a size-based filtering strategy that removes exceptionally small objects lacking distinctive features. This filtering process is essential for maintaining object integrity and preventing the inclusion of ambiguous or noisy visual elements that could hinder model performance.
@@ 3.2 Restoration-based Supervised Fine-Tuning
Fig. 4 illustrates the architecture of our proposed R-SFT framework, a novel training paradigm designed to unify and enhance image generation and editing capabilities within a single model. The framework is structured into three core, interconnected components: (a) Random Mask, (b) Fine-Tuning, and (c) Prompt Alignment, which collectively guide the model from data preparation to high-fidelity output generation. The process initiates with the random mask module (a), which preprocesses the original training images by applying variable-shaped and sized occlusions. This deliberate corruption transforms the task into the restoration problem, forcing the model to learn robust contextual reasoning and inpainting skills essential for both conditional generation and editing. The fine-tuning component (b) forms the core of the training pipeline, implemented as a two-stage progressive learning strategy. In Stage 1, the model is trained to reconstruct images from large-area masks, focusing on recovering broad scene context and global consistency. Stage 2 refines these capabilities using smaller, object-level masks, honing the model’s ability to perform detailed, precise restoration. Both stages leverage a single-stream diffusion transformer  [qin2025lumina] as a backbone, which is efficiently optimized using a combination of LoRA [hu2022lora] and FFN fine-tuning. The details of the DiT architecture are shown in Fig. 5. In the context of diffusion transformer models, a notable advantage of the single-stream architecture lies in its capacity to seamlessly incorporate supplementary conditional inputs. Unlike alternative approaches that necessitate the introduction of additional module branches, such as adapters or ControlNet [zhang2023adding], the single-stream DiT can directly integrate conditional information without the need for such structural modifications. To leverage this characteristic for the inpainting task, we introduce a dedicated conditional embedder. This embedder is specifically designed to process the inpainting image, denoted as $I_{m}$, obtained by element-wise multiplication of the original image $I$ and the mask $m$. The output of the conditional embedder is then concatenated with noisy latent representations and text tokens. Subsequently, a self-attention mechanism is employed to facilitate multi-modal interaction among these concatenated features. This approach enables the model to effectively combine visual information from the inpainting region, latent representations of the corrupted image, and semantic information from the text prompt, thereby enhancing the overall inpainting performance.
@@
Crucially, the Prompt Alignment module (c) operates during inference to ensure semantic consistency. It utilizes a powerful pretrained Large-Language Model (Qwen-3B [bai2023qwen]) to interpret the user’s textual prompt and align it to the caption distribution. This mechanism provides explicit semantic guidance, enabling the model to perform prompt-aware restoration and object edits. By integrating a restoration-based objective with a structured fine-tuning schedule and a dedicated prompt-alignment mechanism, the R-SFT framework effectively equips a unified model with powerful and versatile image manipulation capabilities.
@@ 3.3 Training objective
We employ flow matching to optimize the parameters of our model. In contrast to Denoising Diffusion Probabilistic Models (DDPM) [ho2020denoising], flow matching performs the forward process by linearly interpolating between noise and data along a straight trajectory. At any given timestep $t$, the noisy sample $x_{t}$ is defined as: $x_{t}=t\cdot x+(1-t)\cdot\epsilon$, where x represents the original data and $\epsilon\sim\mathcal{N}(0,I)$ denotes Gaussian noise. The model is trained to directly regress the target velocity vector $v$, conditioned on the noisy input $x_{t}$, the timestep $t$, and any auxiliary conditioning information $c$. The flow matching loss:
@@
Additionally, we introduce a region-weighted loss to direct the model’s focus towards the masked regions that require inpainting or editing. This is a critical need for generating small, distinct objects like cones and pedestrians in traffic scenes. Diffusion models often fail to consider fine-grained details in localized edits. This can result in blurry or inconsistent generation in small areas. This phenomenon arises from the standard mean-squared error loss function, which treats all pixels in an identical manner. Consequently, the gradient signal from a small edited region is overwhelmed by the substantial unchanged background. The weighted loss as follows:
$w_{i,j}=\left\{\begin{array}[]{ll}1&\text{if }\mathbf{I}_{i,j}=\mathbf{I}_{i,j}^{\prime}\\ \frac{1}{\|\mathbf{I}-\mathbf{I}^{\prime}\|^{2}}&\text{if }\mathbf{I}_{i,j}\neq\mathbf{I}_{i,j}^{\prime}\end{array}\right$. (2)
@@ 4.1 Experiment Details
Our dataset consists of three components: open-source datasets (BDD100K [yu2020bdd100k], TT100K [zhu2016traffic], CUHK-SYSU [xiao2017joint]), an internal proprietary dataset, and an internet dataset. The training set comprises approximately 2.4M text-image pairs, including 2M pairs with detailed global image descriptions and 0.4M pairs with local object-level descriptions. The validation set is carefully curated, consisting of 100 samples. This data encompasses a wide variety of scenes in the traffic domain, including different perspectives, elements, weather conditions, and lighting variations. Our model is trained on 8xA800 GPUs using the AdamW [loshchilov2017decoupled] optimizer, with an initial learning rate of 1e-4 and a global batch size of 32. The training supports a dynamic resolution strategy and is conducted in two stages. In the first stage, we train for 16 epochs using global image descriptions to capture the high-level semantic structure of traffic scenes. In the second stage, we fine-tune for 5 epochs, using local object descriptions to refine details and enhance local realism. Evaluation metrics such as FID [heusel2017gans], SSIM, and PSNR are used to assess the model’s convergence and generation quality.
@@ 4.3 Quantative Analysis
We evaluated our two-stage model on the validation set, as shown in Tab. 1. In Stage 1, we pre-trained the model using large-scale image-text pairs, achieving FID = 9.8, SSIM = 0.77, and PSNR = 21.06. In Stage 2, we fine-tuned the model with fine-grained object descriptions, resulting in improved performance: FID decreased to 9.7, while SSIM and PSNR increased to 0.79 and 24.13, respectively. An ablation study confirmed that our masked region-weighted loss function decreased FID further, from 9.7 to 9.6, while PSNR and SSIM increased to 0.81 and 24.52 respectively. This demonstrates that focusing loss weighting on object areas improves both detail fidelity and semantic accuracy.